OpenStudy (anonymous):

(Fan+ Medal) -Please- Indicate which property is illustrated in Step 4. Step 1 2 • 9 • 11 • 1 = (2 • 9) • (11 • 1) Step 2 = (9 • 2) • (1 • 11) Step 3 = 18 • (1 • 11) Step 4 = 18 • 11 A. identity property of multiplication B. identity property of addition C. distributive property D. inverse property of multiplication

OpenStudy (lifeisadangerousgame):

Do you know what each of these properties mean?

OpenStudy (anonymous):

No

OpenStudy (lifeisadangerousgame):

Okay, I'll give you the definition: Identity property of multiplication: This property says that a number does not change when that number is multiplied by 1. (example: 2 * 1 = 1 or 5678493 * 1 = 5678493) Distributive Property: This is a property that is used to multiply a single term and two or more terms inside a set of parentheses. (example: 3(5 + x) = 15 + 3x you multiply 3 by both terms/numbers/variables inside of the parentheses) Inverse Property of Multiplication: This property says that if you multiply a number by its reciprocal (multiplicative inverse) the product is 1. (example: 1/4 * 4 = 1 or 4 * 1/4 = 1) The Identity Property of Addition: This property says that zero added to any number is the number itself. (example: 3 + 0 = 3 or 5 + 0 = 5)

OpenStudy (zaberangelr):

i think a